OpenStudy (anonymous):

determine if the following pairs are equal by changing both to the same denominator 11/15 and 25/33 are the fractions equals? equal or not equal?

OpenStudy (mertsj):

The common denominator is 3(5)(11) = 165 11/15 = 121/165 25/33= 125/165 So no. they are not equal.